A club treasurer has $28.45 consisting of nickles, dimes, and quarters. She has 9 more nickles than dimes and three times as many dimes as quarters. How many does she have of each coin?
She has ____ nickles, ____ dimes, and ____ quarters.

Answer by stanbon(48535)

A club treasurer has $28.45 consisting of nickles, dimes, and quarters. She has 9 more nickles than dimes and three times as many dimes as quarters. How many does she have of each coin?
---------------------------
Let # of quarters be "x"; Value is 25x cents
# of dimes is "3x"; Value is 10*3x = 30x cents
# of nickels is "3x+9" ; Value is 5(3x+9) = 15x+45 cents
--------------------------------
EQUATION:
value + value + value = 2845 cents
25x + 30x + 15x+45 = 2845
70x = 2800
x = 40 (# of quarters)
3x = 120 (# of dimes)
3x+9 = 129 (# of nickels)
=================
Cheers,
Stan H.